# Eliciting Latent Predictions from Transformers with the Tuned Lens

### v5

@@ 1 Introduction
Figure 1: Comparison of our method, the tuned lens (bottom), with the “logit lens” (top) for GPT-Neo-2.7B prompted with an except from the abstract of Vaswani et al. (2017). Each cell shows the top-1 token predicted by the model at the given layer and token index. The logit lens fails to elicit interpretable predictions before layer 21, but our method succeeds.
The impressive performance of transformers in natural language processing (Brown et al., 2020) and computer vision (Dosovitskiy et al., 2020) suggests that their internal representations have rich structure worthy of scientific investigation. One common approach is to train classifiers to extract specific concepts from hidden states, like part-of-speech and syntactic structure (Hewitt and Manning, 2019; Tucker et al., 2021; Li et al., 2022).
In this work, we instead examine transformer representations from the perspective of iterative inference (Jastrzębski et al., 2017). Specifically, we view each layer in a transformer language model as performing an incremental update to a latent prediction of the next token. [^1] See Appendix C for evidence supporting this view, including novel empirical results of our own. We decode these latent predictions through early exiting, converting the hidden state at each intermediate layer into a distribution over the vocabulary. This yields a sequence of distributions we call the prediction trajectory, which exhibits a strong tendency to converge smoothly to the final output distribution, with each successive layer achieving lower perplexity.
We build on the “logit lens” (nostalgebraist, 2020), an early exiting technique that directly decodes hidden states into vocabulary space using the model’s pretrained unembedding matrix. We find the logit lens to be unreliable (Section 2), failing to elicit plausible predictions for models like BLOOM (Scao et al., 2022) and GPT Neo (Black et al., 2021). Even when the logit lens appears to work, its outputs are hard to interpret due to representational drift: features may be represented differently at different layers of the network. Other early exiting procedures also exist (Schuster et al., 2022), but require modifying the training process, and so can’t be used to analyze pretrained models. Simultaneous to this work is (Din et al., 2023) which proposes a relatively similar methodology which we will compare with in future work.
To address the shortcomings of the logit lens, we introduce the tuned lens. We train $L$ affine transformations, one for each layer of the network, with a distillation loss: transform the hidden state so that its image under the unembedding matches the final layer logits as closely as possible. We call these transformations translators because they “translate” representations from the basis used at one layer of the network to the basis expected at the final layer. Composing a translator with the pretrained unembedding yields a probe (Alain and Bengio, 2016) that maps a hidden state to a distribution over the vocabulary.
@@
Finally, we apply the tuned lens in several ways: we extend the results of Halawi et al. (2023) to new models (Section 5.1), we find that tuned lens prediction trajectories can be used to detect prompt injection attacks (Perez and Ribeiro, 2022) often with near-perfect accuracy (Section 5.2), and find that data points which require many training steps to learn also tend to be classified in later layers (Section 5.3).
@@ 2 The Logit Lens
The logit lens was introduced by nostalgebraist (2020), who found that when the hidden states at each layer of GPT-2 (Radford et al., 2019) are decoded with the unembedding matrix, the resulting distributions converge roughly monotonically to the final answer. More recently it has been used by Halawi et al. (2023) to understand how transformers process few-shot demonstrations, and by Dar et al. (2022), Geva et al. (2022), and Millidge and Black (2022) to directly interpret transformer weight matrices.
The method. Consider a pre-LayerNorm transformer [^2] Both the logit lens and the tuned lens are designed primarily for the pre-LN architecture, which is more unambiguously iterative. Luckily pre-LN is by far more common than post-LN among state-of-the-art models. See Zhang and He (2020, Appendix C) for more discussion. $\mathcal{M}$. We’ll decompose $\mathcal{M}$ into two “halves,” $\mathcal{M}_{\leq\ell}$ and $\mathcal{M}_{>\ell}$. The function $\mathcal{M}_{\leq\ell}$ consists of the layers of $\mathcal{M}$ up to and including layer $\ell$, and it maps the input space to hidden states. Conversely, the function $\mathcal{M}_{>\ell}$ consists of the layers of $\mathcal{M}$ after $\ell$, which map hidden states to logits.
@@
While this simple form of the logit lens works reasonably well for GPT-2, nostalgebraist (2021) found that it fails to extract useful information from the GPT-Neo family of models (Black et al., 2021). In an attempt to make the method work better for GPT-Neo, they introduce an extension which retains the last transformer layer, yielding:
@@
Unreliability. Beyond GPT-Neo, the logit lens struggles to elicit predictions from several other models released since its introduction, such as BLOOM (Scao et al., 2022) and OPT 125M (Zhang et al., 2022) (Figure 14).
@@
This is concerning because it suggests we can’t interpret the logit lens prediction trajectory as a belief updating in response to new evidence. The beliefs of a rational agent should not update in an easily predictable direction over time (Yudkowsky, 2007), since predictable updates can be exploited via Dutch books (Ramsey, 1926; De Finetti et al., 1937; Hacking, 1967; Garrabrant et al., 2016). Biased logit lens outputs are trivially exploitable once the direction of bias is known: one could simply “bet” against the logit lens at layer $\ell<L$ that the next token will be one of the tokens that it systematically downweights, and make unbounded profit in expectation.
@@ 3 The Tuned Lens
Representation drift. Another issue with the logit lens is that transformer hidden states often contain a small number of very high variance dimensions, and these “rogue dimensions” (Timkey and van Schijndel, 2021) tend to be distributed unevenly across layers; see Figure 6 (top) for an example. Ablating an outlier direction can drastically harm performance (Kovaleva et al., 2021), so if $\mathrm{LogitLens}$ relies on the presence or absence of particular outlier dimensions, the perplexity of logit lens predictions might be spuriously high.
@@
where $f_{>\ell}(\boldsymbol{h}_{\ell})$ refers to the rest of the transformer after layer $\ell$. This can be viewed as a distillation loss, using the final layer distribution as a soft label (Sanh et al., 2019). It ensures that the probes are not incentivized to learn extra information over and above what the model has learned, which can become a problem when training probes with ground truth labels (Hewitt and Liang, 2019).
Implementation details. When readily available, we train translators on a slice of the validation set used during pretraining, and use a separate slice for evaluation. Since BLOOM and GPT-2 do not have publicly available validation sets, we use the Pile validation set (Gao et al., 2020; Biderman et al., 2022). The OPT validation set is also not publicly available, but a member of the OPT team helped us train a tuned lens on the OPT validation set. Documents are concatenated and split into uniform chunks of length 2048.
We use SGD with Nesterov momentum, with a linear learning rate decay schedule over 250 training steps. We use a base learning rate of 1.0, or 0.25 when keeping the final transformer layer, and clip gradients to a norm of 1. We accumulate gradients as necessary to achieve a total batch size of $2^{18}$ tokens per optimizer step. We initialize all translators to the identity transform, and use a weight decay of $10^{-3}$.
We evaluate all models on a random sample of 16.4M tokens from their respective pretraining validation sets. We leave out the final transformer layer for GPT-2 (Radford et al., 2019), GPT-NeoX-20B (Black et al., 2022), OPT (Zhang et al., 2022), and Pythia (Biderman et al., 2023), and include it for GPT-Neo (Black et al., 2021). We evaluate BLOOM (Scao et al., 2022) under both conditions in Figure 4.
@@
Results. We plot tuned lens perplexity as a function of depth for the Pythia models and GPT-NeoX-20B in Figure˜5333Pythia and GPT-NeoX-20B were trained using the same architecture, data, and codebase (Andonian et al., 2021). While they’re not officially the same model suite, they’re more consistent than the OPT models.; results for other model families can be found in Appendix A.
@@
As an example of the transferability of lenses to fine-tuned models we use Vicuna 13B (Chiang et al., 2023), an open source instruct fine-tuned chat model based on LLaMA 13B (Touvron et al., 2023). In Figure˜12, Appendix˜A, we compare the performance of a lens specifically trained on Vicuna to a tuned lens trained only on LLaMA. Both Tuned Lenses were trained using a subsample of the RedPajama dataset (Together, 2023), an open source replication of LLaMA’s training set. The performance of both lenses was then evaluated on the test set of Anthropic’s Helpful Harmless (Bai et al., 2022) conversation dataset and the RedPajama Dataset. On the RedPajama Dataset we find, at worst, a 0.3 bits per byte increase in KL divergence to the models final output. On the helpful harmless corpus we find no significant difference between the transferred and trained lenses. These results show that fine-tuning minimally effects the representations used by the tuned lens. This opens applications of the tuned lens for monitoring changes in the representations of a module during fine-tuning and minimizes the need for practitioners to train lenses on new fine-tuned models.
Relation to model stitching. The tuned lens can be viewed as a way of “stitching” an intermediate layer directly onto the unembedding, with an affine transform in between to align the representations. The idea of model stitching was introduced by Lenc and Vedaldi (2015), who form a composite model out of two frozen pretrained models $A$ and $B$, by connecting the bottom layers of $A$ to the top layers of $B$. An affine transform suffices to stitch together independently trained models with minimal performance loss (Bansal et al., 2021; Csiszárik et al., 2021). The success of the tuned lens shows that model stitching works for different layers inside a single model as well.
Benefits over traditional probing. Unlike Alain and Bengio (2016), who train early exiting probes for image classifiers, we do not learn a new unembedding for each layer. This is important, since it allows us to shrink the size of each learned matrix from $|\mathcal{V}|\times d$ to $d\times d$, where $|\mathcal{V}|$ ranges from 50K (GPT-2, Pythia) to over 250K (BLOOM). We observe empirically that training a new unembedding matrix requires considerably more training steps and a larger batch size than training a translator, and often converges to a worse perplexity.
@@ 4 Measuring Causal Fidelity
Prior work has argued that interpretability hypotheses should be tested with causal experiments: an interpretation of a neural network should make predictions about what will happen when we intervene on its weights or activations (Olah et al., 2020; Chan et al., 2022). This is especially important for probing techniques, since it’s known that probes can learn to rely on spurious features unrelated to the model’s performance (Hewitt and Liang, 2019; Belinkov, 2022).
@@
    Latent directions that are important to the tuned lens should also be important to the final layer output. Concretely, if the tuned lens relies on a feature [^4] For simplicity we assume the “features as directions” hypothesis (Elhage et al., 2022), which defines a “feature” to be the one-dimensional subspace spanned by a unit vector $\boldsymbol{v}$. $\boldsymbol{v}$ in the residual stream (its output changes significantly when we manipulate $\boldsymbol{v}$) then the model output should also change a lot when we manipulate $\boldsymbol{v}$.
@@ 4.1 Causal basis extraction
To test Property 1, we first need to find the important directions for the tuned lens. Amnesic probing (Elazar et al., 2021) provides one way to do this—it seeks a direction whose erasure maximally degrades a model’s accuracy.
@@ 4.2 Stimulus-response alignment
We can operationalize this with the Aitchison geometry (Aitchison, 1982), which turns the probability simplex into a vector space equipped with an inner product. In order to downweight the influence of rare tokens, we use the weighted Aitchison inner product introduced by Egozcue and Pawlowsky-Glahn (2016), defined as
@@
We propose to use CBE (Section 4.1) to define a “natural” choice for the intervention $g$. Specifically, for each layer $\ell$, we intervene on the subspace spanned by $\ell$’s top 10 causal basis vectors— we’ll call this the “principal subspace”— using a recently proposed method called resampling ablation (Chan et al., 2022).
@@ 5 Applications
Figure 10: For most models and tasks, we find there is a layer at which the tuned lens performance is better than final layer performance under incorrect demonstrations. Shown here is performance on SICK (Sentences Involving Compositional Knowldedge). Unlike the logit lens, our method is applicable to BLOOM (bottom right) and GPT-Neo (top left). Y-axis shows median-calibrated accuracy as used in Halawi et al. (2023)

### v6

@@ 1 Introduction
Figure 1: Comparison of our method, the tuned lens (bottom), with the “logit lens” (top) for GPT-Neo-2.7B prompted with an except from the abstract of vaswani2017attention. Each cell shows the top-1 token predicted by the model at the given layer and token index. The logit lens fails to elicit interpretable predictions before layer 21, but our method succeeds.
The impressive performance of transformers in natural language processing (brown2020language) and computer vision (dosovitskiy2020image) suggests that their internal representations have rich structure worthy of scientific investigation. One common approach is to train classifiers to extract specific concepts from hidden states, like part-of-speech and syntactic structure (hewitt2019structural; tucker2021if; li2022emergent).
In this work, we instead examine transformer representations from the perspective of iterative inference (jastrzkebski2017residual). Specifically, we view each layer in a transformer language model as performing an incremental update to a latent prediction of the next token. [^1] See Appendix C for evidence supporting this view, including novel empirical results of our own. We decode these latent predictions through early exiting, converting the hidden state at each intermediate layer into a distribution over the vocabulary. This yields a sequence of distributions we call the prediction trajectory, which exhibits a strong tendency to converge smoothly to the final output distribution, with each successive layer achieving lower perplexity.
We build on the “logit lens” (nostalgebraist2020logitlens), an early exiting technique that directly decodes hidden states into vocabulary space using the model’s pretrained unembedding matrix. We find the logit lens to be unreliable (Section 2), failing to elicit plausible predictions for models like BLOOM (scao2022bloom) and GPT Neo (black2021gpt). Even when the logit lens appears to work, its outputs are hard to interpret due to representational drift: features may be represented differently at different layers of the network. Other early exiting procedures also exist (schuster2022confident), but require modifying the training process, and so can’t be used to analyze pretrained models.
To address the shortcomings of the logit lens, we introduce the tuned lens. We train $L$ affine transformations, one for each layer of the network, with a distillation loss: transform the hidden state so that its image under the unembedding matches the final layer logits as closely as possible. We call these transformations translators because they “translate” representations from the basis used at one layer of the network to the basis expected at the final layer. Composing a translator with the pretrained unembedding yields a probe (alain2016understanding) that maps a hidden state to a distribution over the vocabulary.
@@
Finally, we apply the tuned lens in several ways: we extend the results of halawi2023overthinking to new models (Section 5.2), we find that tuned lens prediction trajectories can be used to detect prompt injection attacks (perez2022ignore) often with near-perfect accuracy (Section 5.3), and find that data points which require many training steps to learn also tend to be classified in later layers (Section 5.4).
@@ 2 The Logit Lens
The logit lens was introduced by nostalgebraist2020logitlens, who found that when the hidden states at each layer of GPT-2 (radford2019language) are decoded with the unembedding matrix, the resulting distributions converge roughly monotonically to the final answer. More recently it has been used by halawi2023overthinking to understand how transformers process few-shot demonstrations, and by dar2022analyzing, geva2022transformer, and millidge2022svd to directly interpret transformer weight matrices.
The method. Consider a pre-LayerNorm transformer [^2] Both the logit lens and the tuned lens are designed primarily for the pre-LN architecture, which is more unambiguously iterative. Luckily pre-LN is by far more common than post-LN among state-of-the-art models. See zhang2020accelerating for more discussion. $\mathcal{M}$. We’ll decompose $\mathcal{M}$ into two “halves,” $\mathcal{M}_{\leq\ell}$ and $\mathcal{M}_{>\ell}$. The function $\mathcal{M}_{\leq\ell}$ consists of the layers of $\mathcal{M}$ up to and including layer $\ell$, and it maps the input space to hidden states. Conversely, the function $\mathcal{M}_{>\ell}$ consists of the layers of $\mathcal{M}$ after $\ell$, which map hidden states to logits.
@@
While this simple form of the logit lens works reasonably well for GPT-2, nostalgebraist2021logitlens found that it fails to extract useful information from the GPT-Neo family of models (black2021gpt). In an attempt to make the method work better for GPT-Neo, they introduce an extension which retains the last transformer layer, yielding:
@@
Unreliability. Beyond GPT-Neo, the logit lens struggles to elicit predictions from several other models released since its introduction, such as BLOOM (scao2022bloom) and OPT 125M (zhang2022opt) (Figure 14).
@@
This is concerning because it suggests we can’t interpret the logit lens prediction trajectory as a belief updating in response to new evidence. The beliefs of a rational agent should not update in an easily predictable direction over time (yudkowsky2007conservation), since predictable updates can be exploited via Dutch books (ramsey1926truth; de1937foresight; hacking1967slightly; garrabrant2016logical). Biased logit lens outputs are trivially exploitable once the direction of bias is known: one could simply “bet” against the logit lens at layer $\ell<L$ that the next token will be one of the tokens that it systematically downweights, and make unbounded profit in expectation.
@@ 3 The Tuned Lens
Representation drift. Another issue with the logit lens is that transformer hidden states often contain a small number of very high variance dimensions, and these “rogue dimensions” (timkey2021all) tend to be distributed unevenly across layers; see Figure 6 (top) for an example. Ablating an outlier direction can drastically harm performance (kovaleva2021bert), so if $\mathrm{LogitLens}$ relies on the presence or absence of particular outlier dimensions, the perplexity of logit lens predictions might be spuriously high.
@@
where $f_{>\ell}(\boldsymbol{h}_{\ell})$ refers to the rest of the transformer after layer $\ell$. This can be viewed as a distillation loss, using the final layer distribution as a soft label (sanh2019distilbert). It ensures that the probes are not incentivized to learn extra information over and above what the model has learned, which can become a problem when training probes with ground truth labels (hewitt2019designing).
Implementation details. When readily available, we train translators on a slice of the validation set used during pretraining, and use a separate slice for evaluation. Since BLOOM and GPT-2 do not have publicly available validation sets, we use the Pile validation set (gao2020pile; biderman2022datasheet). The OPT validation set is also not publicly available, but a member of the OPT team helped us train a tuned lens on the OPT validation set. Documents are concatenated and split into uniform chunks of length 2048.
We evaluate all models on a random sample of 16.4M tokens from their respective pretraining validation sets. We leave out the final transformer layer for GPT-2 (radford2019language), GPT-NeoX-20B (black2022gpt), OPT (zhang2022opt), and Pythia (biderman2023pythia), and include it for GPT-Neo (black2021gpt). We evaluate BLOOM (scao2022bloom) under both conditions in Figure 4.
@@
In our initial experiments, we used SGD with Nesterov momentum, with a linear learning rate decay schedule over 250 training steps. We used a base learning rate of 1.0, or 0.25 when keeping the final transformer layer, and clip gradients to a norm of 1. We accumulated gradients as necessary to achieve a total batch size of $2^{18}$ tokens per optimizer step. We initialized all translators to the identity transform, and use a weight decay of $10^{-3}$.
Muon optimizer. After the initial publication of this paper, jordan6muon introduced the Muon optimizer. Using a fast orthogonalization algorithm, Muon attempts to maximize the effective rank (roy2007effective) of the update applied to each matrix-valued parameter, thereby counteracting the well-known tendency of neural network gradients to be rank-deficient. Conceptually, Muon is the opposite of LoRA (hu2022lora), a popular parameter-efficient finetuning method that constrains parameter updates to be low rank. liu2025muon find that Muon accelerates the training of modern large language models compared to AdamW.
Strikingly, we find that using Muon instead of SGD for training the tuned lens dramatically accelerates training, allowing us to achieve much lower KL divergence losses than were possible in our initial experiments. Essentially, all of the tuned lenses trained and evaluated in the earlier versions of this paper were severely undertrained. The weight matrices of Muon-trained lenses have Frobenius norms that are often several times larger than SGD-trained ones, showing that they are much further from the logit lens than our original lenses— they are much better “tuned,” as it were. We opted not to re-do all our experiments with tuned lenses trained using Muon, but we encourage practitioners to use Muon in future experiments.
Results. We plot tuned lens perplexity as a function of depth for the Pythia models and GPT-NeoX-20B in Figure˜5333Pythia and GPT-NeoX-20B were trained using the same architecture, data, and codebase (gpt-neox-library). While they’re not officially the same model suite, they’re more consistent than the OPT models.; results for other model families can be found in Appendix A.
@@
As an example of the transferability of lenses to fine-tuned models we use Vicuna 13B (vicuna2023), an open source instruct fine-tuned chat model based on LLaMA 13B (touvron2023llama). In Figure˜13, Appendix˜A, we compare the performance of a lens specifically trained on Vicuna to a tuned lens trained only on LLaMA. Both Tuned Lenses were trained using a subsample of the RedPajama dataset (redpj2023), an open source replication of LLaMA’s training set. The performance of both lenses was then evaluated on the test set of Anthropic’s Helpful Harmless (bai2022training) conversation dataset and the RedPajama Dataset. On the RedPajama Dataset we find, at worst, a 0.3 bits per byte increase in KL divergence to the models final output. On the helpful harmless corpus we find no significant difference between the transferred and trained lenses. These results show that fine-tuning minimally effects the representations used by the tuned lens. This opens applications of the tuned lens for monitoring changes in the representations of a module during fine-tuning and minimizes the need for practitioners to train lenses on new fine-tuned models.
Relation to model stitching. The tuned lens can be viewed as a way of “stitching” an intermediate layer directly onto the unembedding, with an affine transform in between to align the representations. The idea of model stitching was introduced by lenc2015understanding, who form a composite model out of two frozen pretrained models $A$ and $B$, by connecting the bottom layers of $A$ to the top layers of $B$. An affine transform suffices to stitch together independently trained models with minimal performance loss (bansal2021revisiting; csiszarik2021similarity). The success of the tuned lens shows that model stitching works for different layers inside a single model as well.
Benefits over traditional probing. Unlike alain2016understanding, who train early exiting probes for image classifiers, we do not learn a new unembedding for each layer. This is important, since it allows us to shrink the size of each learned matrix from $|\mathcal{V}|\times d$ to $d\times d$, where $|\mathcal{V}|$ ranges from 50K (GPT-2, Pythia) to over 250K (BLOOM). We observe empirically that training a new unembedding matrix requires considerably more training steps and a larger batch size than training a translator, and often converges to a worse perplexity.
@@ 4 Measuring Causal Fidelity
Prior work has argued that interpretability hypotheses should be tested with causal experiments: an interpretation of a neural network should make predictions about what will happen when we intervene on its weights or activations (olah2020zoom; chan2022causal). This is especially important for probing techniques, since it’s known that probes can learn to rely on spurious features unrelated to the model’s performance (hewitt2019designing; belinkov2022probing).
@@
    Latent directions that are important to the tuned lens should also be important to the final layer output. Concretely, if the tuned lens relies on a feature [^4] For simplicity we assume the “features as directions” hypothesis (elhage2022toy), which defines a “feature” to be the one-dimensional subspace spanned by a unit vector $\boldsymbol{v}$. $\boldsymbol{v}$ in the residual stream (its output changes significantly when we manipulate $\boldsymbol{v}$) then the model output should also change a lot when we manipulate $\boldsymbol{v}$.
@@ 4.1 Causal basis extraction
To test Property 1, we first need to find the important directions for the tuned lens. Amnesic probing (elazar2021amnesic) provides one way to do this—it seeks a direction whose erasure maximally degrades a model’s accuracy.
@@ 4.2 Stimulus-response alignment
We can operationalize this with the Aitchison geometry (aitchison1982statistical), which turns the probability simplex into a vector space equipped with an inner product. In order to downweight the influence of rare tokens, we use the weighted Aitchison inner product introduced by egozcue2016changing, defined as
@@
We propose to use CBE (Section 4.1) to define a “natural” choice for the intervention $g$. Specifically, for each layer $\ell$, we intervene on the subspace spanned by $\ell$’s top 10 causal basis vectors— we’ll call this the “principal subspace”— using a recently proposed method called resampling ablation (chan2022causal).
@@
Figure 10: The tuned lens elicits secrets more effectively than the logit lens at early layers, then performs more poorly at later layers. Interestingly, on this task, the “direct” prompts yield lower elicitation accuracy than “standard” ones.
@@ 5 Applications
We start by reproducing and extending two use cases of the logit lens in the literature.
Figure 11: For most models and tasks, we find there is a layer at which the tuned lens performance is better than final layer performance under incorrect demonstrations. Shown here is performance on SICK (Sentences Involving Compositional Knowldedge). Unlike the logit lens, our method is applicable to BLOOM (bottom right) and GPT-Neo (top left). Y-axis shows median-calibrated accuracy as used in halawi2023overthinking
5.1 Eliciting Secret Knowledge from Language Models
cywinski2025eliciting apply the logit lens to uncover knowledge that an AI possesses but does not verbalize. Specifically, they fine-tune a model to generate hints and respond to user queries about a specific “taboo” word, while never verbalizing it directly. They then prompt an “auditor” LLM to guess what the taboo word is, allowing it to ask the taboo model questions in a black-box fashion. They find that the auditor is much more accurate at guessing the taboo word when it is provided with the top 20 highest probability latent predictions from the taboo model according to the logit lens. This is true in two different settings: “standard,” where the prompt asks for a clue about the taboo word, and “direct,” where the prompt attempts to compel the model to explicitly reveal the taboo word.
We hypothesized that the tuned lens might perform better than the logit lens for this task. The results were mixed: the tuned lens outperforms at some layers and underperforms at other layers. Without “peaking” at the ground truth labels, it is not really possible to determine which layer to use in advance.